# Domain-Invariant Representation Learning of Bird Sounds

### v7

@@ Abstract
Passive acoustic monitoring (PAM) is crucial for bioacoustic research, enabling non-invasive species tracking and biodiversity monitoring. Citizen science platforms provide large annotated datasets from focal recordings, where the target species is intentionally recorded. However, PAM requires monitoring in passive soundscapes, creating a domain shift between focal and passive recordings, challenging deep learning models trained on focal recordings. To address domain generalization, we leverage supervised contrastive learning by enforcing domain invariance across same-class examples from different domains. Additionally, we propose ProtoCLR, an alternative to SupCon loss which reduces the computational complexity by comparing examples to class prototypes instead of pairwise comparisons. We conduct few-shot classification based on BIRB, a large-scale bird sound benchmark to assess pre-trained bioacoustic models. Our findings suggest that ProtoCLR is a better alternative to SupCon. [^1] Models and code: https://github.com/ilyassmoummad/ProtoCLR
keywords:
contrastive learning, prototypical learning, domain generalization, few-shot learning, bioacoustics
@@ 1 Introduction
Domain Generalization (DG) [8] aims to develop models that learn robust features that are domain-invariant, i.e. capable of generalizing to new unseen domains without prior knowledge or access to target domain data during training. SupCon offers a promising approach for DG. In SupCon, the objective is to learn an embedding space where same-class examples are pulled together and different-class examples are pushed apart. This clustering can promote domain-invariance when sufficient domain diversity is present in the dataset, allowing the model to focus on features that are domain-invariant. In contrast, its self-supervised counterpart SimCLR [9] lacks this explicit mechanism for domain-invariance, as it relies solely on augmentations to create positive pairs. Without label information, SimCLR requires carefully designed augmentations that account for domain shift [10].
Despite its effectiveness, SupCon is computationally expensive due to the need for pairwise similarity calculations between all examples. To address this, in this paper we introduce ProtoCLR (Prototypical Contrastive Learning of Representations), a more efficient variant. By analyzing SupCon's gradient and drawing inspiration from the generalization capabilities of prototypical learning [11, 12] in few-shot learning, ProtoCLR replaces pairwise comparisons with a prototypical contrastive loss that compares examples to class prototypes (one prototype per class), retaining the original objective while significantly reducing computational complexity.
We propose a new few-shot classification evaluation based on the BIRB [13] benchmark to evaluate the generalization capabilities of models trained on XC's focal recordings and tested on diverse soundscape datasets. This benchmark is designed to assess how well models can generalize across domains in challenging few-shot scenarios. We validate our proposed loss ProtoCLR on this benchmark, demonstrating its effectiveness in improving DG in bird sound classification.
@@ 2 Related Work
DG [8] has emerged as a critical approach to tackle domain shift, where the test data distribution differs from the training data. It aims to learn robust, domain-invariant representations using only source domain data.DG methods typically focus on learning domain-invariant representations, using techniques like domain alignment [18], meta-learning [19], and data augmentation [20]. These approaches help the model learn features that remain consistent across varying domains. In bioacoustics, DG is especially important due to the difficulty in collecting annotated soundscape recordings compared to focal data [13, 17].
Invariant learning has gained attention, where models are trained to learn features that remain invariant across different variations in data, such as augmented versions in self-supervised learning [9, 21, 22, 23] or same-class examples in supervised learning [5]. Moummad et al. [6] showed that invariant learning effectively learns representations of bird sounds that perform well in few-shot learning tasks for recognizing new species not seen during training. However, their experiments focus on 5-way 1-shot classification tasks, that do not represent realistic scenarios.In our study, we examine test datasets featuring passively recorded sounds from a variable number of species, reflecting more realistic and unencountered training scenarios.
@@ 3 Method
Figure 1: Prototypical contrastive pre-training of representations: Anchors are augmented and pulled towards the prototype of their class, while being pushed apart from prototypes of other classes.
@@ 4 Experiments
Table 1: Overview of BIRB datasets.
@@ 4.1 Few-Shot Classification Benchmark
The original task of BIRB [13] is information retrieval in bioacoustics, focused on retrieving bird vocalizations from passively recorded datasets using focal recordings for training. Downstream datasets may contain long audio recordings where events of interest are sparse. Individual bird events are detected using peak detection [24], which identifies frames with high energy to used for the task. We build on previous work [6], which shows that using a pretrained classifier on AudioSet [25] is a good proxy for selecting windows with the highest bird class activation.
@@ 4.2 Reference Systems
To compare supervised contrastive approaches SupCon and ProtoCLR for addressing domain generalization, we train a reference system using cross-entropy (CE) loss as a supervised baseline and SimCLR as a self-supervised contrastive baseline. Additionally, we evaluate large-scale, state-of-the-art models in bioacoustics: BirdAVES [29], a transformer model based on the speech model HuBERT, trained in a self-supervised manner based on the AVES framework [29]. We evaluate two versions of BirdAVES, biox-base, a 12-layer transformer with 768 units, trained on Xeno-Canto [3] and animal sounds from AudioSet [25] and VGGSound [30], and bioxn-large, a 24-layer transformer with 1024 units, trained on the same datasets plus iNaturalist  [^2] https://github.com/gvanhorn38/iNatSounds. BirdAVES outputs one embedding vector per time frame, which we average across the time dimension to produce a single vector representing the entire audio clip; the encoder of BioLingual [31], an HT-SAT transformer pre-trained on AudioSet and fine-tuned using contrastive language-audio training to align animal sounds with text captions describing the class across a large collection of data including Xeno-Canto, iNaturalist, Animal Sound Archive, …etc; and the encoder of Perch [13], a large EfficientNet [32] trained on Xeno-Canto for species classification, as well as taxonomic ranks genus, family, and order. We evaluate only models trained specifically on bird sounds, as general-purpose audio models have been shown to perform poorly in this domain [26]. Similarly, Hamer et al. [13] found that general-purpose models can, in some cases, perform worse than simply averaging mel spectrogram features without any learning.
@@ 4.4 Results and Discussion
Table 2 presents the performance of different models on 1-shot and 5-shot bird sound classification tasks. ProtoCLR outperforms SupCon in both settings, while also being more computationally efficient during training: for 1 epoch on XC with a batch size of 256, SupCon requires 80.4B multiply-accumulate operations, whereas ProtoCLR requires only 28.3B.

### v8

@@ Abstract
Passive acoustic monitoring (PAM) is crucial for bioacoustic research, enabling non-invasive species tracking and biodiversity monitoring. Citizen science platforms provide large annotated datasets from focal recordings, where the target species is intentionally recorded. However, PAM requires monitoring in soundscapes, creating a domain shift between focal and passive recordings, challenging deep learning models trained on focal recordings. To address domain generalization, we leverage supervised contrastive learning by enforcing domain invariance across same-class examples from different domains. Additionally, we propose ProtoCLR, an alternative to SupCon loss which reduces the computational complexity by comparing examples to class prototypes instead of pairwise comparisons. We conduct few-shot classification based on BIRB, a large-scale benchmark to assess pre-trained bioacoustic models. Our findings suggest that ProtoCLR is a better alternative to SupCon. [^1] Code: https://github.com/ilyassmoummad/ProtoCLR.
This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011015097 made by GENCI.
Index Terms—  Contrastive learning, prototypical learning, domain generalization, few-shot learning, bioacoustics.
@@ 1 Introduction
Domain Generalization (DG) [33] aims to develop models that learn robust features that are domain-invariant, i.e. capable of generalizing to new unseen domains without prior knowledge or access to target domain data during training. SupCon offers a promising approach for DG. In SupCon, the objective is to learn an embedding space where same-class examples are pulled together and different-class examples are pushed apart. This clustering can promote domain-invariance when sufficient domain diversity is present in the dataset, allowing the model to focus on features that are domain-invariant, without relying on any domain annotations. In contrast, its self-supervised counterpart SimCLR [5] lacks this explicit mechanism for domain-invariance, as it relies solely on augmentations to create positive pairs. Without label information, SimCLR requires carefully designed augmentations that account for domain shift [32].
Despite its effectiveness, SupCon is computationally expensive due to the need for pairwise similarity calculations between all examples. To address this, in this paper we introduce ProtoCLR (Prototypical Contrastive Learning of Representations), a more efficient variant. By analyzing SupCon’s gradient and drawing inspiration from the generalization capabilities of prototypical learning [24, 16] in few-shot learning, ProtoCLR replaces pairwise comparisons with a prototypical contrastive loss that compares examples to class prototypes (one prototype per class), retaining the original objective while significantly reducing computational complexity.
We propose a new few-shot classification evaluation based on the BIRB [11] benchmark to evaluate the generalization capabilities of models trained on XC’s focal recordings and tested on diverse soundscape datasets. This benchmark is designed to assess how well models can generalize across domains in challenging few-shot scenarios. We validate our proposed loss ProtoCLR on this benchmark, demonstrating its effectiveness in improving DG in bird sound classification.
@@ 2 Related Work
DG [33] has emerged as a critical approach to tackle domain shift, where the test data distribution differs from the training data. It aims to learn robust, domain-invariant representations using only source domain data.DG methods typically focus on learning domain-invariant representations, using techniques like domain alignment, meta-learning and data augmentation [33]. These approaches help the model learn features that remain consistent across varying domains. In bioacoustics, DG is especially important due to the difficulty in collecting annotated soundscape recordings compared to focal data [11, 22].
Invariant learning has gained attention, where models are trained to learn features that remain invariant across different variations in data, such as augmented versions in self-supervised learning [5, 3, 31] or same-class examples in supervised learning [15]. Moummad et al. [18] showed that invariant learning effectively learns representations of bird sounds that perform well in few-shot learning tasks for recognizing new species not seen during training. However, their experiments focus on 5-way 1-shot classification tasks, that do not represent realistic scenarios.In our study, we examine test datasets featuring passively recorded sounds from a variable number of species, reflecting more realistic and unencountered training scenarios.
@@ 3 Method
Fig. 1: Prototypical contrastive pre-training of representations: Anchors are augmented and pulled towards the prototype of their class, while being pushed apart from prototypes of other classes.
@@ 4 Experiments
Table 1: Overview of the datasets in the BIRB benchmark.
@@ 4.1 Few-Shot Classification Benchmark
The original task of the benchmark BIRB [11] focuses on information retrieval in bioacoustics, aiming to retrieve bird vocalizations from passively recorded datasets using focal recordings for training. Downstream datasets may contain long audio recordings where events of interest are sparse. Individual bird events are detected using peak detection [6], which identifies frames with high energy to used for the task. We build on previous work [18], which shows that using a pretrained classifier on AudioSet [8] is a good proxy for selecting windows with the highest bird class activation.
@@ 4.2 Reference Systems
To compare supervised contrastive approaches SupCon and ProtoCLR for addressing domain generalization, we train a reference system using cross-entropy (CE) loss as a supervised baseline and SimCLR as a self-supervised contrastive baseline. Additionally, we evaluate large-scale, state-of-the-art models in bioacoustics: BirdAVES [10], a transformer model based on the speech model HuBERT, trained in a self-supervised manner based on the AVES framework [10]. We evaluate two versions of BirdAVES, biox-base, a 12-layer transformer with 768 units, trained on Xeno-Canto [28] and animal sounds from AudioSet [8] and VGGSound [4], and bioxn-large, a 24-layer transformer with 1024 units, trained on the same datasets plus iNaturalist  [^2] https://github.com/gvanhorn38/iNatSounds. BirdAVES outputs one embedding vector per time frame, which we average across the time dimension to produce a single vector representing the entire audio clip; the encoder of BioLingual [23], an audio transformer pre-trained on AudioSet and fine-tuned using contrastive language-audio training to align animal sounds with text captions describing the class across a large collection of data including Xeno-Canto, iNaturalist, Animal Sound Archive, …etc; and the encoder of Perch [11], a large EfficientNet [27] trained on Xeno-Canto for species classification, as well as taxonomic ranks genus, family, and order. We evaluate only models trained specifically on bird sounds, as general-purpose audio models have been shown to perform poorly in this domain [9]. Similarly, Hamer et al. [11] found that general-purpose models can, in some cases, perform worse than simply averaging mel spectrogram features without any learning.
@@ 4.4 Results and Discussion
Table 2 presents the performance of different models on 1-shot and 5-shot bird sound classification tasks. ProtoCLR outperforms SupCon, while being computationally efficient during training: 1 epoch on XC with a batch size of 256, SupCon requires 80.4B multiply-accumulate operations, whereas ProtoCLR requires only 28.3B.